OpenStudy (anonymous):

For the function f(x) = 2 – 3x, find f(–4).

OpenStudy (anonymous):

I got -10

OpenStudy (anonymous):

double check your calculations f(-4)= 2-3(-4)

OpenStudy (anonymous):

f(-4)=2+12

OpenStudy (anonymous):

f(-4)=14?

OpenStudy (anonymous):

So it is 14?

OpenStudy (jdoe0001):

\(\bf f({\color{red}{ x}}) = 2 - 3{\color{red}{ x}}\qquad \qquad f({\color{red}{ -4}})=2 - 3({\color{red}{ -4}})\)

OpenStudy (anonymous):

f(-4)=14? @jdoe0001

OpenStudy (jdoe0001):

-3*-4 = +12 + 2 = 14 yes

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

Thanks, @jdoe0001 and @d3v13

OpenStudy (jdoe0001):

yw